Please help me solve these problems
3x-5y=-8 & 6x+61=y

Answer by gonzo(654)

a straight substitution should solve this.
two equations.
3*x-5*y = (-8)
6*x+61 = y
substitute 6*x+61 for y in the first equation which becomes
3*x-5*(6*x+61)=(-8) which becomes
3*x-30*x-305=(-8) which becomes
-27*x=297 which becomes
-x=11 which becomes
x=-11
substituting x = -11 in the first equation yields
3*(-11)-(5*y)=(-8)
-33-(5*y)=(-8)
-5*y=25
-y=5 which becomes
y=-5
answer is x = -11 and y = -5
substituting in the original equation we get 3*(-11)-5*(-5)=-8 which becomes
-33+25=-8
which becomes -8=-8 proving values for x and y are correct.
in the second equation 6(-11)+61=-5 which becomes
-66+61=-5 which becomes
-5=-5 proving the values are correct for the second equation.